# The Open Proof Corpus: A Large-Scale Study of LLM-Generated Mathematical Proofs

### v1

@@ 3.3 Monitoring and Validation
To ensure grading quality and judge consistency, we implemented a set of monitoring and validation procedures.
@@
One experienced judge was assigned as coordinator, responsible for tracking grading progress, resolving issues, and ensuring motivation. As a senior author, the coordinator had detailed knowledge of the dataset’s goals and methodology and was available to answer judges’ questions throughout the process.
@@
Prior to full-scale grading, we conducted a test run with a limited number of problems. This helped us identify interface issues, unclear instructions, and other inefficiencies. Based on feedback, we revised the interface and grading guidelines before launching the full evaluation.
@@
Approximately $10\%$ of the proofs were double-graded to evaluate inter-judge consistency. Disagreements were reviewed by the coordinator to determine whether they arose from misunderstandings, ambiguous instructions, or errors overlooked by a judge. If possible, instructions were further improved to prevent similar issues. However, most inconsistencies stemmed from overlooked errors in the proofs and could therefore not be resolved by clarifying the instructions. If the coordinator identified a significant number of discrepancies for a specific judge, they would discuss the issue with the judge to clarify instructions.
@@ 4 Open Proof Corpus
Figure 2: Overview of competitions included in the OPC.
We now introduce the OPC, providing an overview of the competitions and models it includes and key dataset statistics. At a high level, the OPC is built to support the training and evaluation of proof generation models and to facilitate a rigorous analysis of their capabilities.
@@
The OPC consists of 5,0625{,}062 proofs across 1,0101{,}010 distinct problems, generated by six state-of-the-art LLMs. Each proof is labeled as either correct or incorrect by one or two human judges. Labels are accompanied by short justifications, with optional annotations highlighting specific sentences within the proof. Each problem may also include metadata such as its competition source, difficulty level, and other relevant attributes.
@@
Problems were sourced from a wide range of prestigious mathematics competitions. A full breakdown is provided in App.˜A, with a summary shown in Fig.˜2. App.˜A also reports the average accuracy of the best-performing model per competition, offering a rough proxy for difficulty. Most problems are at the high school level, with only a small portion drawn from undergraduate-level competitions. Competitions were selected to cover a broad range of difficulties while ensuring an overall average model accuracy of approximately $50\%$.
Table 1: Number of solutions evaluated per model.
@@
Table˜1 summarizes the number of proofs generated by each model in the OPC. Not all models contributed to every problem, but most problems include solutions from at least five models. o4-mini contributed the largest share of proofs, as it was used extensively in best-of-n and pass@n experiments.
Human performance and noise
To estimate label reliability, we double-graded approximately $10\%$ of the dataset. Among these, judges agreed on the proof’s correctness in $90.4\%$ of cases. Assuming independent judgments, we can estimate the individual judge error rate $p$ by solving $0.904=(1-p)^{2}+p^{2}$, giving $p=5\%$. This indicates a relatively low noise level, which is expected for a human-annotated dataset of this complexity.
@@
  • MathArena: A test set of $112$ problems from MathArena (Balunović et al., 2025), a final-answer benchmark for mathematical reasoning.
@@
  • Best-of-n: A subset of $152$ problems from the IMOSL, USAMO, and BMOSL, each solved multiple times by o4-mini. For 60 of these problems, all 8 generations were human-evaluated. The rest include judgments only for the generation selected by a best-of-n selection strategy.
  • Generic: A subset of $676$ problems from various competitions, including the IMOSL, USAMO, BMOSL, and others, all solved by multiple models.
The MathArena and PutnamBench subsets are drawn from existing public benchmarks and should be treated strictly as test sets. The generic and best-of-n subsets, by contrast, are intended for training, validation, and further analysis. However, a small portion of the generic subset is held back for benchmarking purposes as well.
@@ 5 Results
We present the main findings derived from the OPC. In Section˜5.1, we evaluate the proof-generation capabilities of various models. Section˜5.2 evaluates the ability of LLMs to judge the correctness of proofs. In Section˜5.3, we compare informal and formal proof-generation performance. Section˜5.4 analyzes proof correctness given correct final answers. Finally, Section˜5.5 examines the effectiveness of best-of-n selection strategies. In App.˜C, we additionally provide several qualitative observations about common mistakes in the generated proofs.
5.1 Gemini-2.5-Pro Is The Best Proof Generator
Figure 3: Average proof correctness of various models on the OPC. Data is split into two partitions, the first, resp. second, containing only problems answered by all models except DeepSeek-R1, resp. Grok-3-Mini. The discrepancy between the two partitions is due to the inclusion of harder competitions in the second partition.
Fig.˜3 shows the average proof correctness of each model on a filtered subset of the OPC, including only questions answered by all models except Grok-3-Mini or DeepSeek-R1. Gemini-2.5-Pro achieves the highest average accuracy, slightly outperforming o4-mini. In contrast, the two open-source models—Qwen3-235B-A22B and DeepSeek-R1—underperform significantly, highlighting the current performance gap between closed-source and open-source models in proof generation
Out of more than 1,700 incorrect solutions analyzed, models explicitly state their inability to solve the problem in only 114 instances, with all but five of those generated by o3. Even o3 is more likely to produce an incorrect proof than to acknowledge uncertainty. This widespread reluctance to admit failure highlights a key limitation: current LLMs lack effective mechanisms for confidence calibration and knowing their limitations. In domains like mathematics, this could undermine trust in systems that rely on LLMs for provably correct solutions, especially given the difficulty of evaluating proofs correctly.
5.2 LLMs Are Near-Human Level Judges
Table 2: Benchmarking LLMs as proof graders. Cost for running the model on the entire subset is given in USD.
@@ 5.2 LLMs Are Near-Human Level Judges
Baseline $53.2$ $53.2$ N/A
The OPC is fundamentally a dataset of binary human judgments on proof correctness, making it well-suited for training and evaluating LLMs as proof judges. To leverage this, we split the generic subset by problem into training and test sets. Using GRPO (Shao et al., 2024), we fine-tune DeepSeek-R1-Qwen3-8B on the training set with human labels for reward modeling.
The test set comprises 293 LLM-generated proofs. We evaluate both reasoning models, such as Gemini-2.5-Pro and o4-mini, and general-purpose models like GPT-4.1. As a baseline, we classify proofs under 100 characters as incorrect and guess randomly otherwise.
As shown in Table˜2, Gemini-2.5-Pro achieves the highest judging accuracy: $85.4\%$ with a single evaluation pass and $88.1\%$ with majority voting, approaching the $90.4\%$ human baseline. Notably, OPC-R1-8B matches Gemini-2.5-Pro’s majority voting performance and outperforms its base model by $17\%$, demonstrating the value of the OPC and its potential for advancing the field of proof evaluation and generation. However, since the train and test sets share the same distribution, OPC-R1-8B’s performance may degrade on out-of-distribution data.
Table 3: Judgement accuracy breakdown split by solver, highlighting the lowest score for each judge. For visual clarity, we shortened Gemini-2.5-Pro to Gemini, o4-mini to o4, and Qwen3-235B-A22B to Qwen.
Judge Prover Gemini o4 o3 Qwen
Gemini-2.5-Pro 79.4 $87.1$ $91.6$ $80.6$
o4-mini $86.9$ 81.3 $83.1$ $84.1$
o3 $85.9$ $84.8$ 76.9 $87.8$
Qwen3-235B-A22B $80.0$ $81.9$ 79.1 $84.4$
LLMs often favor their own generations (Panickssery et al., 2024). To investigate this, we evaluate how well models judge their proofs compared to those generated by others. In Table˜3, we find that all models except Qwen3-235B-A22B perform significantly worse when judging their own proofs, indicating an inability to self-critique effectively. This suggests that LLMs struggle to recognize their mistakes, which is a critical limitation for applications requiring self-assessment or iterative improvement.
This positive result appears to contradict with Petrov et al. (2025), who reported poor performance of judge models. However, their evaluation was based on a limited set of problems, relied on older models, and focused on the more challenging task of scoring proofs on a continuous rather than a binary scale. In addition, we put considerable effort in crafting clear and comprehensible prompt instructions.
@@ 5.3 Formal Proof Generation Lags Behind
Using the PutnamBench subset, we compare formal and natural language proof-generation models. The best formal model, DeepSeek-Prover-V2 (Ren et al., 2025), achieves less than $8\%$ accuracy on PutnamBench. In contrast, Fig.˜4 shows that the top informal model, Gemini-2.5-Pro, reaches almost $83\%$ accuracy on the evaluated subset, clearly outperforming all formal models. Despite this disparity, formal proofs offer a major advantage: automatic verifiability. While informal methods currently dominate in performance, formal approaches remain crucial for scalable, rigorous proof checking.
@@ 5.4 Proof Generation and Final Answer Do Not Align
Although it is widely claimed that final-answer benchmarks are inadequate for evaluating proof generation capabilities (Petrov et al., 2025; Mahdavi et al., 2025; Guan et al., 2025), it remains unclear how often LLMs can produce a valid proof when they find the correct answer. To investigate this, we first collect instances from the MathArena subset where models generate correct final answers, and then manually evaluate the validity of the accompanying proofs. This enables us to estimate the overall proof correctness rate P(correct proof)P($\text${correct proof}) and compare it with the final-answer accuracy P(correct final answer)P($\text${correct final answer}).
In Fig.˜5, we report each model’s accuracy alongside the stricter metric requiring both a correct final answer and a valid proof. Despite Gemini-2.5-Pro, o4-mini, and o3 achieving similar final-answer accuracy, their proof correctness rates differ significantly. Specifically, o3 performs notably worse, with only $59.5\%$ of its answers containing a correct proof. This substantial variation across models highlights that final-answer accuracy alone is not a reliable indicator of proof generation capability.
@@ 5.5 Best-of-n Significantly Improves Performance
Figure 6: Pass@n metric on a small subset of the OPC compared with some best-of-n selection strategies.
@@
Best-of-n selection—generating multiple outputs and selecting the best one—is a common strategy for improving LLM performance. We evaluate this approach using o4-mini by generating eight proofs per problem in the best-of-n subset and testing four selection methods:
@@
  • Discrete: o4-mini classifies each proof as correct or incorrect and selects any correct proof.
@@
  • Continuous: o4-mini scores proofs on a 0-77 scale and selects the one with the highest score.
@@
  • Rank (Bracket): A ranking method proposed by Liu et al. (2025). Proofs are judged pairwise by o4-mini in a knockout tournament until one proof remains. Note that Liu et al. (2025) finetuned their model on the judging task, while we use o4-mini itself.
@@
  • Rank (Swiss): Inspired by Liu et al. (2025), proofs are ranked via pairwise comparisons in a Swiss round-robin tournament. Ratings are computed using the Bradley-Terry model (Bradley and Terry, 1952), and the answer with the highest rating is selected. See App.˜B for details of this method.
@@
Prompts for all methods can be found in App.˜E. Note that Rank (Bracket) only requires $O(n)$ comparisons, and is therefore as efficient as the discrete and continuous methods, while Rank (Swiss) requires $O(n^{2})$ comparisons, making it more expensive.
@@
In Fig.˜6, we compare the performance of these methods with the pass@n metric on the $60$ problems that contain human judgments for all eight proofs. We find that best-of-n selection strategies can significantly improve proof generation performance. Furthermore, the pairwise ranking methods significantly outperform the discrete and continuous methods by approximately $10\%$. Notably, while discrete and continuous methods plateau after $n=5$, ranking approaches continue to scale. Finally, Rank (Swiss) slightly outperforms Rank (Bracket) by a $3\%$ margin.
Figure 7: Best-of-n selection strategies on the entire best-of-n subset of the OPC.
@@
In Fig.˜7, we evaluate the performance of the best-of-n selection strategies on the entire best-of-n subset of the OPC, except for the Rank (Bracket) method, which was not evaluated on the full subset. The improved performance of the ranking methods is confirmed on the entire best-of-n subset, with Rank (Swiss) improving accuracy by $17\%$.

### v2

@@ 3.3 Monitoring and Validation
To ensure grading quality and judge consistency, we implemented a set of validation procedures.
@@
One experienced judge was assigned as coordinator, responsible for tracking grading progress, resolving issues, and ensuring motivation. As a core author, the coordinator had detailed knowledge of the dataset’s goals and methodology and was available to answer judges’ questions.
Developer
In addition to the coordinator, one of the core authors served as developer, providing technical support for the grading interface and addressing any software-related issues that arose during the grading process. Judges were encouraged to report bugs or suggest improvements to the grading interface, which the developer would address immediately to ensure a smooth grading experience. The developer was also responsible for implementing new features, such as the LLM-generated issue summaries, and to monitor model performance metrics to inform problem selection and judge assignments.
@@
Before full-scale grading, we conducted a test run with a limited number of problems. In particular, four experienced judges graded approximately 300 proofs in this initial stage. During this phase, a more significant portion of proofs (around $35\%$) were double-graded to evaluate inter-judge consistency. Judges were also very actively monitored, and encouraged to actively ask questions about any small ambiguity they encountered. After this stage, judge feedback was collected to improve the grading interface and instructions. In particular, the instructions were refined to address remaining ambiguities and edge cases.
@@
In total, approximately $10\%$ of the proofs were double-graded to evaluate inter-judge consistency. Throughout the grading pipeline, disagreements were reviewed by the coordinator to determine whether they arose from misunderstandings, ambiguous instructions, or errors overlooked by a judge. If possible, instructions were further improved to prevent similar issues. However, most inconsistencies came from overlooked errors in the proofs and could therefore not be resolved by clarifying the instructions. If the coordinator identified a significant number of discrepancies for a specific judge, they would discuss the issue with the judge to clarify instructions.
@@ 4 Open Proof Corpus
We now introduce the OPC and provide an overview of key dataset statistics. At a high level, the OPC is built to support training and to facilitate a rigorous analysis of proof generation capabilities.
Figure 2: Overview of competitions in the OPC.
@@
The OPC consists of 5,0625{,}062 proofs across 1,0101{,}010 distinct problems, generated by six state-of-the-art LLMs. Each proof is labeled as either correct or incorrect by one or two human judges. Labels are accompanied by short justifications, with optional annotations highlighting specific sentences within the proof. Each problem may also include metadata such as its competition source, difficulty level, and other relevant attributes. An example is shown in Fig.˜1(a).
@@
Problems were sourced from a wide range of prestigious mathematics competitions. A full breakdown is provided in App.˜A, with a summary shown in Fig.˜2. App.˜A also reports the average accuracy of the best-performing model per competition, offering a rough proxy for difficulty. Most problems are at the high school level ($\approx 84\%$), with a small portion drawn from undergraduate-level competitions ($\approx 16\%$).
Table 1: Solutions per evaluated model.
@@
Table˜1 shows the number of proofs generated by each model. Not all models attempted every problem, but most problems include solutions from five models. o4-mini contributed the largest number of proofs, as it was used in the best-of-n and pass@n experiments.
Human performance
To estimate label reliability, we double-graded approximately $10\%$ of the dataset. Among these, judges agreed on the proof’s correctness in $90.4\%$ of cases. Assuming independent judgments, we can estimate the individual judge error rate $p$ by solving $0.904=(1-p)^{2}+p^{2}$, giving $p=5\%$. This indicates a low noise level, which is expected for a human-annotated dataset of this complexity.
@@
  • MathArena: A subset of $112$ problems from MathArena (Balunović et al., 2025), a final-answer benchmark for mathematical reasoning. $34$ questions from the SMT 2025 will only be released after the competition authors have published the questions publicly.
@@
  • Best-of-n: A subset of $152$ problems from hard competitions, each solved multiple times by o4-mini. For 60 of these problems, all 8 generations were human-evaluated. The rest include judgments only for the generations selected by a best-of-n selection strategy.
  • Generic: A subset of $676$ problems from various competitions, solved by multiple models.
The MathArena and PutnamBench subsets are drawn from existing benchmarks and should not be used for training. The generic and best-of-n subsets are intended for training, validation, and further analysis. However, a small portion of the generic subset is held back for benchmarking purposes.
@@ 5 Results
We now present our main findings. In Section˜5.1, we evaluate the proof-generation capabilities of various models. Section˜5.2 evaluates the ability of LLMs to judge proof correctness. In Section˜5.3, we compare informal and formal proof-generation performance. Section˜5.4 analyzes proof correctness given correct final answers. Section˜5.5 examines the effectiveness of various best-of-n strategies. Finally, in Section˜5.6, we discuss the potential impact of data contamination on our results. In App.˜E, we additionally provide some qualitative observations about common mistakes in the generated proofs.
All shown confidence intervals in this section are 95% intervals computed using the large sample normal approximation. Bold numbers indicate the best performance in the respective category, and underlined numbers are within the confidence interval of the best performance.
5.1 Proof Generation Capabilities
Figure 3: Average proof correctness on the OPC. Data is split into two partitions, the first, resp. second, containing only problems answered by all models except R1, resp. Grok-3-Mini. The second partition contains problems from more challenging competitions, explaining the score discrepancy.
Proof generation results
Fig.˜3 shows the average proof correctness of each model on a subset of the OPC, including only fully-judged questions for all models except Grok-3-Mini or R1. Gemini-2.5-Pro achieves the highest accuracy, slightly outperforming o4-mini. In contrast, the two open-source models, Qwen3-235B-A22B and R1, underperform significantly, highlighting the performance gap between closed- and open-source models in generating correct proofs.
Uncertainty acknowledgment
Out of more than 1,700 incorrect solutions analyzed, models explicitly state their inability to solve the problem in only 114 instances, with all but five of those generated by o3. Even o3 is more likely to produce an incorrect proof than to acknowledge uncertainty. This reluctance to admit failure highlights a key limitation. In domains like mathematics, this could undermine trust in systems that rely on LLMs for provably correct solutions.
5.2 LLMs Are Human Level Judges
Setup
The OPC provides binary human judgments for proof correctness, making it ideal for training and evaluating LLM proof judges. To leverage this, we split the generic subset by problem into training and test sets. Using GRPO (Shao et al., 2024), we fine-tune R1-Qwen3-8B using human labels for rewards. The test set comprises 293 LLM-generated proofs with 40% average correctness. We evaluate both reasoning models, such as GPT-5, and general-purpose models like GPT-4.1. Importantly, the human baseline is not measured on the test subset, but rather on all double-graded proofs in the OPC. Since the test samples are uniformly drawn from the OPC, this does not significantly affect the comparison.
Table 2: LLMs as proof graders. Cost for running the model on the entire subset is given in USD. Full confidence intervals are given in Section˜C.4.
@@ 5.2 LLMs Are Human Level Judges
Results
As shown in Table˜2, GPT-5 achieves the highest judging accuracy: $89.3\%$ with a single evaluation pass, approaching the $90.4\%$ human baseline, and $90.8\%$ with majority voting. Notably, OPC-R1-8B matches Gemini-2.5-Pro’s majority voting performance and outperforms its base model by $17\%$, demonstrating the value of the OPC and its potential for advancing the field of proof evaluation and generation. However, the train set for OPC-R1-8B shares the same distribution as this test set, which may inflate its performance. In App.˜C, we show that while the performance of OPC-R1-8B is reduced under out-of-distribution data, the improvement over the base model persists even under these conditions.
This positive result appears to contradict Petrov et al. (2025), who reported poor performance of judge models. However, their evaluation was based on a limited set of problems, relied on older models, and focused on the more challenging task of scoring proofs on a continuous rather than a binary scale. In addition, we put considerable effort into crafting clear and comprehensible prompts.
Table 3: Judgement accuracy breakdown, highlighting the lowest score for each judge. Full confidence intervals are given in Section˜C.4.
Prover Judge Gemini o4 o3 Qwen
Gemini 79.4 $86.9$ $85.9$ 80.0
o4 $87.1$ 81.3 $84.8$ 81.9
o3 $91.6$ 83.1 76.9 79.1
Qwen 80.6 84.1 $87.8$ $84.4$
Self-evaluation
LLMs often favor their own generations (Panickssery et al., 2024). To investigate this, we evaluate how well models judge their proofs compared to those generated by others. In Table˜3, we find that all models except Qwen3-235B-A22B perform worse when judging their own proofs. This suggests that LLMs struggle to recognize their own mistakes, which is a critical limitation for applications requiring self-evaluation.
@@ 5.3 Formal Proof Generation Lags Behind
Using the PutnamBench subset, we compare formal and natural language proof-generation models. The best formal model, Goedel-Prover-v2 (Lin et al., 2025b), achieves less than $19\%$ accuracy on the PutnamBench. In contrast, Fig.˜5 shows that the top informal model, Gemini-2.5-Pro, reaches almost $83\%$ accuracy on the evaluated subset, clearly outperforming this baseline. Despite this disparity, formal proofs offer a major advantage: automatic verifiability. While informal methods currently dominate, formal approaches remain crucial for scalable proof verification.
Finally, we note that a recent private agentic system, Seed-Prover (Chen et al., 2025), obtained a $50\%$ formal accuracy on the PutnamBench. However, our informal results do not use agentic techniques, and it is therefore not accurate to directly compare these numbers.
@@ 5.4 Proof Generation and Final Answer Do Not Align
Although it is widely claimed that final-answer benchmarks are inadequate for evaluating proof generation capabilities (Petrov et al., 2025; Mahdavi et al., 2025; Guan et al., 2025), it remains unclear how often LLMs can produce a valid proof when they find the correct answer. To investigate this, we first collect instances from the MathArena subset where models generate correct final answers, and then manually evaluate the validity of the accompanying proofs. This enables us to estimate the overall proof correctness rate and compare it with the final-answer accuracy.
Results
In Fig.˜5, we report each model’s final-answer accuracy alongside the stricter metric requiring a valid proof. Despite Gemini-2.5-Pro, o4-mini, and o3 achieving similar final-answer accuracy, proof correctness differs significantly. Specifically, o3 performs notably worse, with only $59.5\%$ of its answers containing a correct proof. This substantial difference between models shows that final-answer accuracy is not a reliable indicator of proof generation capability.
@@ 5.5 Best-of-n Significantly Improves Performance
((a)) Pass@n versus all other best-of-n methods.
((b)) Selection of methods on a larger subset.
Figure 6: Performance of the best-of-n selection strategies.
Best-of-n selection strategies
@@
Best-of-n selection, generating multiple outputs and selecting the best one, is a common strategy to improve performance. We evaluate this approach using o4-mini by generating eight proofs per problem in the best-of-n subset and testing four selection methods:
@@
  • Discrete: o4-mini classifies each proof as correct or incorrect and selects any correct one.
@@
  • Continuous: o4-mini scores proofs on a 0-77 scale and selects one with the highest score.
@@
  • Rank (Bracket): A ranking method proposed by Liu et al. (2025). Proofs are judged pairwise by o4-mini in a knockout tournament until one proof remains.
@@
  • Rank (Swiss): Inspired by Liu et al. (2025), proofs are paired in a Swiss round-robin tournament. Ratings are computed using the Bradley-Terry model (Bradley and Terry, 1952), and the proof with the highest rating is selected. See App.˜B for details.
Prompts for all methods can be found in App.˜I.
Complexity
@@
Note that Rank (Bracket) requires $O(n)$ comparisons, making it as efficient as the discrete method, while Rank (Swiss) requires $O(n^{2})$ comparisons, making it more expensive.
Results on small subset
@@
In Fig.˜6(a), we compare the performance of these methods with the pass@n metric on the $60$ problems that contain human judgments for all eight proofs. We find that best-of-n selection strategies can significantly improve proof generation performance. Furthermore, the pairwise ranking methods outperform the discrete and continuous methods by approximately $10\%$. Notably, while discrete and continuous methods plateau after $n=5$, ranking approaches continue to scale. Finally, Rank (Swiss) slightly outperforms Rank (Bracket) by a $3\%$ margin.
Results on larger subset
@@
In Fig.˜6(b), we evaluate the performance of the best-of-n selection strategies on the $134$ problems of the best-of-n subset of the OPC [^1] A small bug in the Rank (Swiss) method caused incorrect selections for 18 questions. These are excluded from the analysis., except for the Rank (Bracket) method, which was not evaluated on the full subset. The improved performance of the ranking methods is confirmed on the entire best-of-n subset, with Rank (Swiss) improving accuracy by $17\%$. While the confidence intervals are relatively large due to the small number of problems in this problem set, all selection methods rely on the same underlying answers from o4-mini, making the relative performance differences significant.
5.6 Potential Impact of Contamination
Since many OPC problems are publicly available, there is a possibility that some were contaminated during training. In this section, we briefly explain why this plays an insignificant role in our results.
Proof generation capabilities
In App.˜C, we present a small experiment indicating that contamination has a smaller effect than problem difficulty. Nevertheless, it cannot be fully ruled out. In particular, the small gap in performance between Gemini-2.5-Pro and o4-mini in Section˜5.1 cannot be conclusively attributed to genuine performance differences. However, results on smaller, contamination-free datasets (Petrov et al., 2025; Balunović et al., 2025) also show Gemini-2.5-Pro outperforming o4-mini, supporting the interpretation that the performance gap is real.
In all other sections, the conclusions are robust to contamination: the informal-formal gap is too large to be affected by small changes, the discrepancy between final-answer accuracy and proof correctness is based on problems from MathArena, a well-known benchmark with problems created in 2025, and best-of-n strategies are compared using the same model, making contamination irrelevant.
Table 4: Model accuracy with and without providing the ground truth solutions. $\Delta$ shows the change in accuracy between the two settings.
Judge pass@1 pass@1 (w/ Solution) $\Delta$
GPT-5 89.3 89.0 -0.3
Gemini-2.5-Pro 85.4 82.7 -2.7
OPC-R1-8B 83.8 83.1 -0.7
o4-mini 83.8 85.2 +1.4
o3 83.1 85.9 +2.8
Qwen3-235B-A22B 81.8 84.9 +3.1
R1 80.9 80.8 -0.1
R1-Qwen3-8B 70.7 75.4 +4.7
Proof judging capabilities
Data contamination poses a less significant risk for proof judging, since generated proofs cannot be present in the training data. The underlying problems and their official solutions, however, may have been seen during training, which could give models an advantage. To test this, we run a worst-case experiment where the ground-truth solution is provided alongside the proof to be judged. As shown in Table˜4, the resulting accuracy gains are small and non-significant, suggesting that access to the solution does little to improve error detection and that contamination has a limited impact on judging performance.